OpenStudy (cindy):

5(2-j) + (2j-3) 10-5j +2j-3 10-3j-3 = 7-3j is that correc t

OpenStudy (immanuelv):

yes ... your work is correct...and so is the answer....!

OpenStudy (anonymous):

correct cindy

OpenStudy (cindy):

is this ne correct can you guys check please

OpenStudy (immanuelv):

its not.... :(

OpenStudy (immanuelv):

13 is not divisible by 4

OpenStudy (cindy):

its actually a 12 not a 13

OpenStudy (immanuelv):

the correct answer is (12/13) - (6x/13) this is the final answer! :)

OpenStudy (immanuelv):

does it make sense?

OpenStudy (cindy):

still dont understand sorry

OpenStudy (immanuelv):

none of them are divisible..... therefore your only hope is to multiply them out when you multiply (1/13)(12-6x) together, you get 12/13 - 6x/13 now does it at least makes sum kinda sense?

OpenStudy (cindy):

sorry its actually 1/3 not 1/13

OpenStudy (immanuelv):

well then it is 4-2x

OpenStudy (cindy):

how canyou show me the steps

OpenStudy (immanuelv):

of course... so your steps were correct but u forgot about the 6. when you multiply (1/3)(12-6x) together, you get (1/3)(12) - (1/3)(-6x) basically you are dividing them by 3 12 /3 =4 -6x/3= -2x therefore it is 4-2x

OpenStudy (immanuelv):

do you understand?

OpenStudy (cindy):

iis equation correct

OpenStudy (immanuelv):

if it is (3-4 y ) then yes!!!! :)

OpenStudy (cindy):

wat do u mean

OpenStudy (immanuelv):

i thought i saw a "y" but never mind .... it is -1

OpenStudy (cindy):

so my equation is correct

OpenStudy (immanuelv):

.... what happened to the y on the top?